### v2

# On kinks and solitons in discrete nonlinear transmission line

Abstract
We studied lossless discrete transmission lines constructed from ideal linear inductors and nonlinear capacitors. The travelling waves with finite support in such systems are the kinks and the solitons. We calculated the speeds and the profiles both of the kinks and of the solitons. We also studied dissipative kinks in the lossy transmission lines, with ohmic resistors shunting the inductors. The losses influence the dissipative kink speed and profile.
I Introduction
The nonlinear electrical transmission lines are of much interest both due to their applications, and as the laboratories to study nonlinear waves malomed2 . Among the latter – the travelling waves, especially the kinks and the solitons, attract a lot of attention solitons . The nonlinear waves become especially interesting (and challenging) when one takes into account the discrete nature of the transmission line. The kinks and the solitons in discrete transmission lines were studied in quite a few publications malomed4 ; cameroon ; bogdan .
The field of nonlinear waves in discrete transmission lines is closely connected with Fermi-Pasta-Ulam-Tsingou (FPUT) problem fermi . In particular, dynamics of lattice kinks and solitons were studied in a large number of publications kevrekidis ; vainchtein
We studied previously the kinks and the solitons in discrete Josephson transmission line (JTL) JTL kogan2 ; kogan3 . There we formulated the quasi-static approximation, which allows to reduce the discrete equations to the continuous ones. In the present short note we would like to apply this approximation to different kind of discrete transmission lines – with nonlinear capacitors. Such transmission lines in the continuum approximation were studied in the classical papers by R. Landauer landauer1 ; landauer2 ; landauer .
II The kinks and the solitons
II.1 The travelling waves
The discrete transmission line is constructed from the identical inductors and nonlinear capacitors is shown on Fig. 1.
Figure 1: Discrete transmission line.
We take the voltages $V_{n}$ and the currents which pass though the inductors $I_{n}$ as the dynamical variables. The circuit equations are
$\frac{dQ_{n}}{dt}$ $=I^{c}_{n},$ (1a)
$L\frac{dI^{c}_{n}}{dt}$ $=V_{n+1}-2V_{n}+V_{n-1},$ (1b)
where $Q_{n}=Q(V_{n})$ and the nonlinear nature of the transmission line is determined by the function $Q(V)$. We also took into account the connection between the current through the capacitor $dI^{c}_{n}$ and the currents between the inductors
$I^{c}_{n}=I_{n-1,n}-I_{n,n+1}.$ (2)
Let us write down Eq. (II.1) in the quasi-continuum approximation, which was proposed in our previous publications kogan2 ; kogan3 . The approximation consists in treating $n$ as the continuous variable $Z$, expanding the r.h.s. of the equation (1b) in Taylor series with respect to $Z$ and keeping only the first two terms of the expansion
$V_{n+1}-2V_{n}+V_{n-1}=\frac{\partial^{2}V}{\partial Z^{2}}+\frac{1}{12}\frac{\partial^{4}V}{\partial Z^{4}}.$ (3)
After that (II.1) takes the form
$\frac{\partial Q}{\partial t}$ $=I^{c},$ (4a)
$L\frac{\partial I^{c}}{\partial t}$ $=-\frac{\partial}{\partial Z}\left(\frac{\partial V}{\partial Z}+\frac{1}{12}\frac{\partial^{3}V}{\partial Z^{3}}\right).$ (4b)
In this paper we are interested in the travelling wave solutions of (II.1), that is the dependence of all the variables upon $Z$ and $t$ is of the form
$X(Z,t)=X(Z-Ut),$ (5)
where $U$ is the speed of the travelling wave (the distance is measured in the units of the transmission line spacial period $\Lambda$). This gives us the opportunity to perform integration in (4b) to get
$ULI^{c}=\frac{\partial V}{\partial Z}+\frac{1}{12}\frac{\partial^{3}V}{\partial Z^{3}}.$ (6)
Substituting (6) into (4a) we obtain ordinary differential equation
$LU^{2}\frac{dQ}{dZ}=\frac{dV}{dZ}+\frac{1}{12}\frac{d^{3}V}{dZ^{3}}.$ (7)
An obvious way to integrate (7) starts by going to the second order ODE
$LU^{2}Q=V+\frac{1}{12}\frac{d^{2}V}{dZ^{2}}+\text{const}.$ (8)
However in this paper we’ll go along the different way. Presenting $dQ/dz$ as $dQ/dz=(dQ/dV)(dV/dZ)$, introducing new dependent variable
$E=\frac{dV}{dZ}$ (9)
and considering $V$ as the independent variable, we may after taking into account that
$\frac{d^{2}V}{dZ^{2}}$ $=\frac{1}{2}\frac{dE^{2}}{dV},$ (10a)
$\frac{d^{3}V}{dZ^{3}}$ $=\frac{1}{2}E\frac{d^{2}E^{2}}{dV^{2}}$ (10b)
and dividing both parts by the common multiplier $E$ in both parts, reduce the order of (7) and write it down as
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}=LU^{2}\frac{dQ(V)}{dV}-1.$ (11)
A simple mechanical interpretation of Eq. (11) may be helpful. If we call $V$ the ”time”, the equation can be considered as describing motion of the fictitious Newtonian particle, $E^{2}$ being its coordinate and $1/24$ it’s mass, under the influence of the time-dependent force, given by the r.h.s. of (11).
We’ll consider travelling waves with finite support, that is impose upon $V(Z)$ the boundary conditions
$\lim_{Z\to-\infty}V(Z)=V_{1},\hskip 28.45274pt\lim_{Z\to+\infty}V(Z)=V_{2}.$ (12)
The value of $V_{2}$ may be different from $V_{1}$, in which case we’ll talk about the kink, or equal to $V_{1}$, in which case we’ll talk about the soliton.
The quantity $E^{2}(V)$ obviously has minimum (equal to zero) at $V=V_{1,2}$. Hence from (11) follows
$LU^{2}\left(\frac{dQ}{dV}\right)_{V=V_{1,2}}-1>0.$ (13)
The inequality (13) has a simple physical meaning. Consider the ”sound” waves that is small amplitude, smooth waves propagating on the homogeneous backgrounds $V=V_{1,2}$. These waves are described by the equation obtained by linearizing (7) near $V=V_{1,2}$ and ignoring the term with the third order derivative
$Lu^{2}\left(\frac{dQ}{dV}\right)_{V=V_{1,2}}\frac{dV}{dZ}=\frac{dV}{dZ}.$ (14)
Hence the speed of the sound is
$\frac{1}{Lu_{1,2}^{2}}=\left(\frac{dQ}{dV}\right)_{V=V_{1,2}}.$ (15)
Now we understand that (13) means
$U^{2}>u_{1,2}^{2},$ (16)
that is the travelling waves are supersonic kogan2 ; kogan3 .
The boundary conditions (12) obviously give
$E\left(V_{1,2}\right)=0.$ (17)
From (10a) follows
$\left(\frac{dE^{2}}{dV}\right)_{V=V_{1,2}}=0.$ (18)
The conditions (17) and (18) mean that the particle is at rest at the origin at ”time” equal to $V_{1}$ and then again at ”time” equal to $V_{2}$.
In the vicinity of $V_{1,2}$ Eq. (11) has the form
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}=LU^{2}\left(\frac{dQ}{dV}\right)_{V=V_{1,2}}-1=\text{const}.$ (19)
Taking into account the boundary conditions (17) and (18) we obtain
$E^{2}\sim\left(V-V_{1,2}\right)^{2},\hskip 28.45274pt\text{ in the vicinity of }V_{1,2}.$ (20)
If we approximate $Q(V)$ between $V_{1}$ and $V_{2}$ by the third degree polynomial
$Q(V)=\text{const}+C_{0}\left(\alpha V+\beta V^{2}+\gamma V^{3}\right),$ (21)
Eq. (11) can be specified to
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}=LU^{2}(\alpha+2\beta V+3\gamma V^{2})-1.$ (22)
For brevity we’ll call Eq. (21) and (22) the cubic case.
II.2 The kinks
With four boundary conditions on $E(V)$ supplied by (17) and (18) for the kink problem of the three parameters $V_{1},V_{2},U^{2}$ only one (say $V_{1}$) can be considered as a free parameter, the other two being connected with it. One such connection we could obtain at much earlier stage. In fact, integrating (7) with respect to $V$ from $V_{1}$ to $V_{2}$ and taking into account the boundary conditions we obtain the speed of the kink
$\frac{1}{LU^{2}}=\frac{Q_{2}-Q_{1}}{V_{2}-V_{1}}.$ (23)
where $Q_{i}=Q(V_{i})$.
If $Q(V)$ is a polynomial function of $V$, from (20) follows that the kink can be presented as
$E^{2}(V)=P_{1}(V)(V_{1}-V)^{2}(V-V_{2})^{2}.$ (24)
In the cubic case, calculating the integral in (23) we obtain
$\frac{1}{LC_{0}U^{2}}=\alpha+\beta\left(V_{1}+V_{2}\right)+\gamma\left(V_{1}^{2}+V_{1}V_{2}+V_{2}^{2}\right).$ (25)
We also can specify (24) to
$E^{2}(V)=\psi^{2}(V_{1}-V)^{2}(V-V_{2})^{2},$ (26)
where $\psi$ is a constant. Substituting (26) into (22) and using Vieta’s formula we obtain after simple algebra
$\psi^{2}$ $=6U^{2}LC_{0}\gamma,$ (27a)
$V_{1}+V_{2}$ $=-\frac{2\beta}{3\gamma}.$ (27b)
To find $V(Z)$ we have to solve equation
$\frac{dV}{dZ}=E(V).$ (28)
The solution is
$V(Z)=\frac{V_{2}+V_{1}}{2}+\frac{V_{2}-V_{1}}{2}\tanh\left[\frac{V_{2}-V_{1}}{2}\psi Z\right]$ (29)
(to satisfy the boundary conditions we have to choose positive value of $\psi$). After finding $V(Z)$ we can from (1b) obtain the current
$I(Z)=\frac{\psi\left(V_{2}-V_{1}\right)}{2LU}\tanh\left[\frac{V_{2}-V_{1}}{2}\psi Z\right].$ (30)
II.3 The solitons
To specify the soliton we need to introduce an additional parameter, and as such we’ll take the extremal value of the voltage $V_{3}$. That is $V$ changes between $V_{1}$ and $V_{3}$, and $|V_{3}-V_{1}|$ we’ll call the amplitude of the soliton. Because $V_{3}$ is the extremal value of $V$ we have
$\left(\frac{dV}{dZ}\right)_{V=V_{3}}=0\Rightarrow E(V_{3})=0.$ (31)
Hence
$E^{2}\sim\left(V-V_{3}\right),\hskip 28.45274pt\text{ in the vicinity of }V_{3}.$ (32)
Recalling our mechanical analogy we see that the soliton corresponds to the motion of the fictitious particle with zero initial coordinate and velocity at $V=V_{1}$ and zero coordinate (but not velocity) at $V=V_{3}$. Mechanical interpretation of (11) in the case of the soliton leads to the situation which doesn’t normally occur in mechanics. The boundary condition $V_{2}=V_{1}$ means that the ”time” $V$ flows along the contour from $V_{1}$ to $V_{3}$ and then back from $V_{3}$ to $V_{1}$. This is why $E(V)$ for the soliton will later turn out to be a double-valued function.
With three boundary conditions on $E(V)$ supplied by (17), (18) and (31) for the soliton problem of the three parameters $V_{1}V_{3},U^{2}$ only two (say $V_{1}$ and $V_{3}$) can be considered as free parameters. Multiplying both sides of (7) by $V_{3}-V$, integrating the equation with respect to $V$ from $V_{1}$ to $V_{3}$ and taking into account the boundary conditions we obtain the speed of the soliton
[…]
If $Q(V)$ is a polynomial function of $V$, from (20) and (32) follows that the soliton can be presented as
$E^{2}(V)=P_{2}(V)(V_{1}-V)^{2}(V-V_{3}),$ (34)
where $P_{1}(V)$ and $P_{2}(V)$ are some polynomial functions.
In the cubic case, calculating the integral in (33) we obtain
$\frac{1}{LC_{0}U^{2}}=\alpha+\frac{2\beta}{3}\left(2V_{1}+V_{3}\right)+\frac{\gamma}{2}\left(3V_{1}^{2}+2V_{1}V_{3}+V_{3}^{2}\right).$
We also can specify (34) (generalize (26)) to
$E^{2}(V)=\psi(V_{1}-V)^{2}(V-V_{3})(V-V_{4}),$ (36)
where $\psi$ is a constant. Substituting (26) into (22) and using Vieta’s formula we obtain after simple algebra
$\psi^{2}=6U^{2}LC_{0}\gamma,$ (37a)
$2V_{1}+V_{3}+V_{4}=-\frac{4\beta}{3\gamma}.$ (37b)
To make life easier to ourselves, let us restrict the consideration by the case $|\beta/\gamma|\gg|V_{1}|,|V_{3}|$. Physically the approximation means that $V_{1}$ and $V_{3}$ are close enough to each other, and we can ignore the cubic term in (21). In this case, excluding $V_{4}$, we obtain that (36) and (II.3) are modified to
$E^{2}(V)$ $=\phi(V_{1}-V)^{2}(V-V_{3}),$ (38a)
$\frac{\phi}{4}$ $=2U^{2}LC_{0}\beta,$ (38b)
$\frac{1}{LC_{0}U^{2}}$ $=\alpha+\frac{2\beta}{3}\left(2V_{1}+V_{3}\right).$ (38c)
From (II.3) we see that the soliton exists both for positive and negative $\beta$. For positive $\beta$ we have dark soliton ($V_{3}<V_{1}$), and for negative $\beta$ – bright soliton ($V_{3}>V_{1}$). From (38c) we see that for any sign of $\beta$ the soliton speed is larger than the sound speed at the background, given by equation
$\frac{1}{LC_{0}u^{2}}=\alpha+2\beta V_{1},$ (39)
in line with our previous general statement that the soliton is supersonic.
To find $V(Z)$ we have to solve equation
$\frac{dV}{dZ}=E(V).$ (40)
The solution of (40) is
$V(Z)=V_{1}+\left(V_{3}-V_{1}\right)\operatorname{sech}^{2}\left(\frac{V_{3}-V_{1}}{2}\phi Z\right).$ (41)
Similar to what was done in the previous subsection, we can use (1b) to find the current and obtain
$I(Z)=\frac{\phi\left(V_{3}-V_{1}\right)}{2LU}\operatorname{sech}^{2}\left(\frac{V_{3}-V_{1}}{2}\phi Z\right).$ (42)
II.4 The quasi-continuum approximation
We would like to discuss an opportunity (and motivation) to go beyond the limits of the quasi-continuum approximation used in the paper. To do it, one may keep the third term in the Taylor expansion of the discrete second order derivative in the r.h.s. of (1b) kogan2 ; kogan3
$V_{n+1}-2V_{n}+V_{n-1}=\frac{\partial^{2}V}{\partial Z^{2}}+\frac{1}{12}\frac{\partial^{4}V}{\partial Z^{4}}+\frac{1}{360}\frac{d^{6}V}{dZ^{6}}.$ (43)
Thus we obtain instead of (7) the equation
$LU^{2}\frac{dQ}{dZ}=\frac{dV}{dZ}+\frac{1}{12}\frac{d^{3}V}{dZ^{3}}+\frac{1}{360}\frac{d^{5}V}{dZ^{5}}.$ (44)
Looking at the results of the previous Sections we understand that the quasi-continuum approximation can be justified if $\gamma(V_{1}-V_{2})^{2}/\alpha\ll 1$ and $\beta\left(V_{1}-V_{3}\right)/\alpha\ll 1$.
The small numerical coefficient before the sixth order derivative in the r.h.s. of (44) means that the small parameters shouldn’t be actually very small. With all that, one can expect that Eq. (44) will allow to study the cases of stronger nonlinearity than Eq. (7).
III The dissipative kinks
Let us ask ourselves what happens with the kinks when the transmission line becomes lossy? To answer this question let us include in the system the resistors $R$ shunting the inductors (see Fig. 2).
Figure 2: Discrete transmission line with shunted inductors.
In this case (2) is modified to
$I^{c}_{n}=I_{n-1,n}-I_{n,n+1}+\frac{1}{R}\left(V_{n+1}-2V_{n}+V_{n-1}\right),$ (45)
hence (II.1) takes the form
$\frac{\partial Q}{\partial t}$ $=I^{c},$ (46a)
$L\frac{\partial I^{c}}{\partial t}$ $=-\frac{\partial}{\partial Z}\left(\frac{\partial V}{\partial Z}+\frac{1}{12}\frac{\partial^{3}V}{\partial Z^{3}}+\frac{L}{R}\frac{\partial^{2}V}{\partial t\partial Z}\right).$ (46b)
(we assume that both the higher order derivatives with respect to $Z$ and $1/R$ are small, so the discrete second order derivative with respect to $Z$ in the r.h.s. of (45) we presented as the continuous second order derivative).
After repeating exactly what was done in Section II.1 we obtain in a modified form Eq. (11)
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+\frac{LU}{R}\frac{dE}{dV}=LU^{2}\frac{dQ(V)}{dV}-1$ (47)
and (in the cubic case) Eq. (22)
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+\frac{LU}{R}\frac{dE}{dV}=LU^{2}(\alpha+2\beta V+3\gamma V^{2})-1.$ (48)
We again impose on $V(Z)$ the boundary conditions (12), which lead to the boundary conditions (17) and (18) for $E(V)$. Equation (47) in the vicinity of $V_{1,2}$ has the form
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+\frac{LU}{R}\frac{dE}{dV}=\text{const},$ (49)
which contains additional term in the l.h.s. with respect to (19), but we still obtain the condition (20). We can also repeat verbatim the part of Section II.2 leading to Eqs. (23) and (24).
In the cubic case we regain Eq. (25), and there exists kogan4 a simple polynomial solution of (48) in the form (26), which we repeat here
$E(V)=\psi(V_{1}-V)(V-V_{2}),$ (50)
where $\psi$ is a constant. Substituting (50) into (48) and equating the coefficients before the same powers of $V$ we obtain after simple algebra the generalization of (II.2)
$\psi^{2}$ $=6U^{2}LC_{0}\gamma,$ (51a)
$V_{1}+V_{2}$ $=-\frac{2\beta}{3\gamma}+\frac{4LU}{\psi R}..$ (51b)
Note that Eqs. (27a), (37a) and (51a) are all identical, hence all the waves we consider in this paper can exist only if $\gamma>0$. Note also that by integrating (48) we can obtain Abel equation of the second kind, and (51b) is the condition of the solvability of the equation polyanin .
Equation (50) shows that in the lossy line the kinks are of the same shape as in the lossless line (given by (29)). However, the value of $V_{2}$ (for a given $V_{1}$) depends upon $R$ (Eq. (51b) instead of (27b)) and so does the kink speed, in spite of the validity in both cases of Eq. (25).
There exists different from (50) type of solutions of (47) (hence different from (29) type of solutions of (III), similar to the shock waves we considered previously kogan1 ; kogan2 ; kogan3 ; kogan4 . In the present paper we don’t consider such solutions.
IV Conclusions
We considered finite support travelling waves of the voltage and the current in the discrete lossless transmission line with nonlinear capacitors. We have shown that these waves are of two kinds: the kinks, when the voltage has different asymptotic values at both sides of the transmission line, and the solitons (solitary waves), when the voltage has the same asymptotic value at both sides of the transmission line. We also studied the dissipative kinks in the lossy discrete transmission line. The losses influence the kink speed and profile.
Acknowledgements.
I am grateful to B. Malomed for the discussion. The paper was finalized during the Author’ participation in Workshop on Twistronics and Moire Materials: Bridging Theory and Experiments, ICTP, Trieste. The Author is grateful to the Center for the hospitality.