### v3

# On the kinks, the solitons and the shocks in discrete nonlinear transmission line

Abstract
We studied lossless discrete transmission lines constructed from ideal linear inductors and nonlinear capacitors. The travelling waves with finite support in such systems are the kinks and the solitons, which speeds and profiles were calculated. We also studied lossy transmission lines. The travelling waves with finite support in such systems are the dissipative kinks and the shocks, which speeds and profiles were also calculated.
I Introduction
The nonlinear electrical transmission lines are of much interest both due to their applications, and as the laboratories to study nonlinear waves malomed2 . Among the latter – the travelling waves, especially the kinks and the solitons, attract a lot of attention solitons . The nonlinear waves become especially interesting (and challenging) when one takes into account the discrete nature of the transmission line. The kinks and the solitons in discrete transmission lines were studied in quite a few publications malomed4 ; cameroon ; bogdan .
The field of nonlinear waves in discrete transmission lines is closely connected with Fermi-Pasta-Ulam-Tsingou (FPUT) problem fermi . In particular, dynamics of lattice kinks and solitons were studied in a large number of publications kevrekidis ; vainchtein
We studied previously the kinks and the solitons in discrete Josephson transmission line (JTL) JTL kogan2 ; kogan3 . There we formulated the quasi-static approximation, which allows to reduce the discrete equations to the continuous ones. In the present short note we would like to apply this approximation to different kind of discrete transmission lines – with nonlinear capacitors. Such transmission lines in the continuum approximation were studied in the classical papers by R. Landauer landauer1 ; landauer2 ; landauer .
II The kinks and the solitons
II.1 The travelling waves
The discrete transmission line is constructed from the identical inductors and nonlinear capacitors is shown on Fig. 1.
Figure 1: Discrete transmission line.
We take the voltages $V_{n}$ and the currents which pass though the inductors $I_{n}$ as the dynamical variables. The circuit equations (Kirchhoff laws) are
$\frac{dQ_{n}}{dt}$ $=I^{c}_{n},$ (1a)
$L\frac{dI_{n,n+1}}{dt}$ $=V_{n}-V_{n+1},$ (1b)
$I^{c}_{n}$ $=I_{n-1,n}-I_{n,n+1}.$ (1c)
where $Q_{n}=Q(V_{n})$ and the nonlinear nature of the transmission line is determined by the (monotonically increasing) function $Q=Q(V)$. We can exclude $I_{n,n+1}$ from (1b) and (1c) and present these two equations a single one
$L\frac{dI^{c}_{n}}{dt}=V_{n+1}-2V_{n}+V_{n-1}.$ (2)
Let us write down Eq. (2) in the quasi-continuum approximation, which was proposed in our previous publications kogan2 ; kogan3 . The approximation consists in treating $n$ as the continuous variable $Z$, expanding the r.h.s. of the equation (2) in Taylor series with respect to $Z$ and keeping only the first two terms of the expansion, thus obtaining the equation in the form
$L\frac{\partial I^{c}_{n}}{\partial t}=\frac{\partial^{2}V}{\partial Z^{2}}+\frac{1}{12}\frac{\partial^{4}V}{\partial Z^{4}}.$ (3)
(And, of course, in the framework of the approximation the derivative in (1a) should be understood as partial derivative.)
From (1a) and (3) we can obtain closed wave equation for the voltage $V$
$L\frac{\partial^{2}Q(V)}{\partial t^{2}}=\frac{\partial^{2}V}{\partial Z^{2}}+\frac{1}{12}\frac{\partial^{4}V}{\partial Z^{4}}.$ (4)
However, further on we will not use Eq. (4); it will be more convenient to use Eqs. (1a) and (3).
In this paper we are interested in the travelling wave solutions, for which the dependence of all the variables upon $Z$ and $t$ is of the form
$X(Z,t)=X(Z-Ut),$ (5)
where $U$ is the speed of the travelling wave (the distance is measured in the units of the transmission line spacial period $\Lambda$). This gives us the opportunity to perform integration in (3) to get
$ULI^{c}=-\frac{\partial V}{\partial Z}-\frac{1}{12}\frac{\partial^{3}V}{\partial Z^{3}}.$ (6)
Substituting (6) into (1a) we obtain ordinary differential equation
$U^{2}L\frac{dQ}{dZ}=\frac{dV}{dZ}+\frac{1}{12}\frac{d^{3}V}{dZ^{3}}.$ (7)
Presenting $dQ/dz$ as $dQ/dz=(dQ/dV)(dV/dZ)$, introducing new dependent variable
$E=\frac{dV}{dZ}$ (8)
and considering $V$ as the independent variable, we may after taking into account that
$\frac{d^{2}V}{dZ^{2}}$ $=\frac{1}{2}\frac{dE^{2}}{dV},$ (9a)
$\frac{d^{3}V}{dZ^{3}}$ $=\frac{1}{2}E\frac{d^{2}E^{2}}{dV^{2}}$ (9b)
and dividing both parts by the common multiplier $E$ in both parts, reduce the order of (7) and write it down as
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}=U^{2}L\frac{dQ(V)}{dV}-1.$ (10)
A simple mechanical interpretation of Eq. (10) may be helpful. If we call $V$ the ”time”, the equation can be considered as describing motion of the fictitious Newtonian particle, $E^{2}$ being its coordinate and $1/24$ it’s mass, under the influence of the time-dependent force, given by the r.h.s. of (10).
We’ll consider travelling waves with finite support, that is impose upon $V(Z)$ the boundary conditions
$\lim_{Z\to-\infty}V(Z)=V_{1},\hskip 28.45274pt\lim_{Z\to+\infty}V(Z)=V_{2}.$ (11)
The value of $V_{2}$ may be different from $V_{1}$, in which case we’ll talk about the kink, or equal to $V_{1}$, in which case we’ll talk about the soliton.
The quantity $E^{2}(V)$ obviously has minimum (equal to zero) at $V=V_{1,2}$. Hence from (10) follows
$U^{2}L\left(\frac{dQ}{dV}\right)_{V=V_{1,2}}-1>0.$ (12)
The inequality (12) has a simple physical meaning. Consider the ”sound” waves that is small amplitude, smooth waves propagating on the homogeneous backgrounds $V=V_{1,2}$. These waves are described by Eq. (10), with the l.h.s. being put equal to zero. Hence the speed of the sound is given by the equation
$\frac{1}{u^{2}L}=\frac{dQ}{dV}.$ (13)
Now we understand that (12) means
$U^{2}>u^{2}\left(V_{1,2}\right),$ (14)
that is the travelling waves are supersonic kogan2 ; kogan3 .
The boundary conditions (11) obviously give
$E\left(V_{1,2}\right)=0.$ (15)
The conditions (15) mean that the particle is at rest at the origin at ”time” equal to $V_{1}$ and then again at ”time” equal to $V_{2}$.
In the vicinity of $V_{1,2}$ Eq. (10) has the form
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}=U^{2}L\left(\frac{dQ}{dV}\right)_{V=V_{1,2}}-1=\text{const}.$ (16)
Taking into account the boundary conditions (15) we obtain
$E\sim\left(V-V_{1,2}\right),\hskip 28.45274pt\text{ in the vicinity of }V_{1,2}.$ (17)
If we approximate $Q(V)$ between $V_{1}$ and $V_{2}$ by the third degree polynomial
$Q(V)=\text{const}+C_{0}\left(\alpha V+\beta V^{2}+\gamma V^{3}\right),$ (18)
Eq. (10) can be specified to
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}=U^{2}L(\alpha+2\beta V+3\gamma V^{2})-1.$ (19)
For brevity we’ll call Eq. (18) and (19) the cubic case. Not that in this case
$\frac{1}{LC_{0}u^{2}(V)}=\alpha+2\beta V+3\gamma V^{2}.$ (20)
II.2 The kinks
With four boundary conditions on $E(V)$ supplied by (15) and (17) for the kink problem of the three parameters $V_{1},V_{2},U^{2}$ only one (say $V_{1}$) can be considered as a free parameter, the other two being connected with it. One such connection we could obtain at much earlier stage. In fact, integrating (7) with respect to $V$ from $V_{1}$ to $V_{2}$ and taking into account the boundary conditions we obtain the speed of the kink
$\frac{1}{U^{2}L}=\frac{Q_{2}-Q_{1}}{V_{2}-V_{1}}.$ (21)
where $Q_{i}=Q(V_{i})$.
If $Q(V)$ is a polynomial function of $V$, from (17) follows that the kink can be presented as
$E^{2}(V)=P_{1}(V)(V-V_{1})^{2}(V-V_{2})^{2}.$ (22)
In the cubic case, calculating the integral in (21) we obtain
$\frac{1}{LC_{0}U^{2}}=\alpha+\beta\left(V_{1}+V_{2}\right)+\gamma\left(V_{1}^{2}+V_{1}V_{2}+V_{2}^{2}\right).$ (23)
We also can specify (22) to
$E^{2}(V)=\psi^{2}(V-V_{1})^{2}(V-V_{2})^{2},$ (24)
where $\psi$ is a constant. Substituting (24) into (19) and using Vieta’s formula we obtain after simple algebra
$\psi^{2}$ $=6U^{2}LC_{0}\gamma,$ (25a)
$V_{1}+V_{2}$ $=-\frac{2\beta}{3\gamma}.$ (25b)
To find $V(Z)$ we have to solve equation
$\frac{dV}{dZ}=E(V).$ (26)
The solution is
$V(Z)=\frac{V_{2}+V_{1}}{2}+\frac{V_{2}-V_{1}}{2}\tanh\left[\frac{V_{2}-V_{1}}{2}\psi Z\right].$ (27)
After finding $V(Z)$ we can from (2) obtain the current
$I(Z)=\frac{\psi\left(V_{2}-V_{1}\right)}{2UL}\tanh\left[\frac{V_{2}-V_{1}}{2}\psi Z\right].$ (28)
II.3 The solitons
To specify the soliton we need to introduce an additional parameter, and as such we’ll take the extremal value of the voltage $V_{3}$. That is $V$ changes between $V_{1}$ and $V_{3}$, and $|V_{3}-V_{1}|$ we’ll call the amplitude of the soliton. Because $V_{3}$ is the extremal value of $V$ we have
$\left(\frac{dV}{dZ}\right)_{V=V_{3}}=0\Rightarrow E(V_{3})=0.$ (29)
Hence
$E^{2}\sim\left(V-V_{3}\right),\hskip 28.45274pt\text{ in the vicinity of }V_{3}.$ (30)
Recalling our mechanical analogy we see that the soliton corresponds to the motion of the fictitious particle with zero initial coordinate and velocity at $V=V_{1}$ and zero coordinate (but not velocity) at $V=V_{3}$. Mechanical interpretation of (10) in the case of the soliton leads to the situation which doesn’t normally occur in mechanics. The boundary condition $V_{2}=V_{1}$ means that the ”time” $V$ flows along the contour from $V_{1}$ to $V_{3}$ and then back from $V_{3}$ to $V_{1}$. This is why $E(V)$ for the soliton will later turn out to be a double-valued function.
With three boundary conditions on $E(V)$ supplied by (15) and (29) for the soliton problem of the three parameters $V_{1},V_{3},U^{2}$ only two (say $V_{1}$ and $V_{3}$) can be considered as free parameters. Multiplying both sides of (7) by $V_{3}-V$, integrating the equation with respect to $V$ from $V_{1}$ to $V_{3}$ and taking into account the boundary conditions we obtain the speed of the soliton
[…]
If $Q(V)$ is a polynomial function of $V$, from (17) and (30) follows that the soliton can be presented as
$E^{2}(V)=P_{2}(V)(V-V_{1})^{2}(V-V_{3}),$ (32)
where $P_{1}(V)$ and $P_{2}(V)$ are some polynomial functions.
In the cubic case, calculating the integral in (31) we obtain
$\frac{1}{LC_{0}U^{2}}=\alpha+\frac{2\beta}{3}\left(2V_{1}+V_{3}\right)+\frac{\gamma}{2}\left(3V_{1}^{2}+2V_{1}V_{3}+V_{3}^{2}\right).$
We also can specify (32) (generalize (24)) to
$E^{2}(V)=\psi(V-V_{1})^{2}(V-V_{3})(V-V_{4}),$ (34)
where $\psi$ is a constant. Substituting (24) into (19) and using Vieta’s formula we obtain after simple algebra
$\psi^{2}=6U^{2}LC_{0}\gamma,$ (35a)
$2V_{1}+V_{3}+V_{4}=-\frac{4\beta}{3\gamma}.$ (35b)
To make life easier to ourselves, let us restrict the consideration by the case $|\beta/\gamma|\gg|V_{1}|,|V_{3}|$. Physically the approximation means that $V_{1}$ and $V_{3}$ are close enough to each other, and we can ignore the cubic term in (18). In this case, excluding $V_{4}$, we obtain that (34) and (II.3) are modified to
$E^{2}(V)$ $=\phi(V-V_{1})^{2}(V-V_{3}),$ (36a)
$\frac{\phi}{4}$ $=2U^{2}LC_{0}\beta,$ (36b)
$\frac{1}{LC_{0}U^{2}}$ $=\alpha+\frac{2\beta}{3}\left(2V_{1}+V_{3}\right).$ (36c)
From (II.3) we see that the soliton exists both for positive and negative $\beta$. For positive $\beta$ we have dark soliton ($V_{3}<V_{1}$), and for negative $\beta$ – bright soliton ($V_{3}>V_{1}$). From (36c) we see that for any sign of $\beta$ the soliton speed is larger than the sound speed at the background, given by equation (20) with $\gamma=0$, in line with our previous general statement that the soliton is supersonic.
To find $V(Z)$ we have to solve equation
$\frac{dV}{dZ}=E(V).$ (37)
The solution of (37) is
$V(Z)=V_{1}+\left(V_{3}-V_{1}\right)\operatorname{sech}^{2}\left(\frac{V_{3}-V_{1}}{2}\phi Z\right).$ (38)
Similar to what was done in the previous subsection, we can use (2) to find the current and obtain
$I(Z)=\frac{\phi\left(V_{3}-V_{1}\right)}{2UL}\operatorname{sech}^{2}\left(\frac{V_{3}-V_{1}}{2}\phi Z\right).$ (39)
II.4 The quasi-continuum approximation
We would like to discuss an opportunity (and motivation) to go beyond the limits of the quasi-continuum approximation used in the paper. To do it, one may keep the third term in the Taylor expansion of the discrete second order derivative in the r.h.s. of (2) kogan2 ; kogan3
$V_{n+1}-2V_{n}+V_{n-1}=\frac{\partial^{2}V}{\partial Z^{2}}+\frac{1}{12}\frac{\partial^{4}V}{\partial Z^{4}}+\frac{1}{360}\frac{d^{6}V}{dZ^{6}}.$ (40)
Thus we obtain instead of (7) the equation
$U^{2}L\frac{dQ}{dZ}=\frac{dV}{dZ}+\frac{1}{12}\frac{d^{3}V}{dZ^{3}}+\frac{1}{360}\frac{d^{5}V}{dZ^{5}}.$ (41)
Looking at the results of the previous Sections we understand that the quasi-continuum approximation can be justified if $\gamma(V_{1}-V_{2})^{2}/\alpha\ll 1$ and $\beta\left(V_{1}-V_{3}\right)/\alpha\ll 1$.
The small numerical coefficient before the sixth order derivative in the r.h.s. of (41) means that the small parameters shouldn’t be actually very small. With all that, one can expect that Eq. (41) will allow to study the cases of stronger nonlinearity than Eq. (7).
III The dissipative kinks and shocks
Let us ask ourselves what happens with the kinks when the transmission line becomes lossy? To answer this question let us include in the system the resistors $R$ shunting the inductors (see Fig. 2).
Figure 2: Discrete transmission line with shunted inductors.
In this case (1c) is modified to
$I^{c}_{n}=I_{n-1,n}-I_{n,n+1}+\frac{1}{R}\left(V_{n+1}-2V_{n}+V_{n-1}\right),$ (42)
hence Eq. (3) takes the form
$L\frac{\partial I^{c}}{\partial t}=\frac{\partial^{2}V}{\partial Z^{2}}+\frac{1}{12}\frac{\partial^{4}V}{\partial Z^{4}}-\frac{L}{R}\frac{\partial^{3}V}{\partial t\partial Z^{2}}.$ (43)
(we assume that both the higher order derivatives with respect to $Z$ and $1/R$ are small, so the discrete second order derivative with respect to $Z$ in the r.h.s. of (42) we presented as the continuous second order derivative).
From (1a) and (43) we can obtain Eq. (4) modified due to the presence of ohmic resistors
$L\frac{\partial^{2}Q(V)}{\partial t^{2}}=\frac{\partial^{2}V}{\partial Z^{2}}+\frac{L}{R}\frac{\partial^{3}V}{\partial t\partial^{2}Z}+\frac{1}{12}\frac{\partial^{4}V}{\partial Z^{4}}.$ (44)
For the travelling wave, combining (1a) and (43), we obtain after integration closed equation for $V(Z)$
$\frac{1}{12}\frac{d^{2}V}{dZ^{2}}+\frac{UL}{R}\frac{dV}{dZ}=U^{2}LQ-V+F,$ (45)
where $F$ is the constant of integration.
Equation (45) is Newton equation, describing motion with friction of the fictitious particle ($V$ being the coordinate of the particle and $Z$ the ”time”) in the force field given by the r.h.s. of the equation. We again impose on $V(Z)$ the boundary conditions (11). Hence the motion starts at $Z=-\infty$ at $V=V_{1}$ and ends at $Z=+\infty$ at $V=V_{2}$ (obviously that both $V_{1}$ and $V_{2}$ are the equilibrium positions). Equating the r.h.s. of (45) to zero at $V=V_{1}$ and $V=V_{2}$ we recover Eq. (21) for the speed of the wave.
It will be convenient to present Eq. (45) as
$\frac{1}{12}\frac{d^{2}V}{dZ^{2}}+\frac{UL}{R}\frac{dV}{dZ}=-\frac{d\Pi(V)}{dV},$ (46)
where the ”potential energy” $\Pi(V)$ is defined by the equation (which took into account the boundary conditions)
$\frac{d\Pi(V)}{dV}=-\frac{V_{2}-V_{1}}{Q_{2}-Q_{1}}Q+V+\frac{V_{2}Q_{1}-V_{1}Q_{2}}{Q_{2}-Q_{1}}.$ (47)
The constraint for the travelling wave asymptotic boundary conditions is the inequality
$\Pi(V_{1})>\Pi(V_{2}).$ (48)
Inequality (48) can be explicitly presented as
[…]
When both $\beta>0$ and $\gamma>0$, the condition (49) means just $|V_{1}|>|V_{2}|$.
The point $V=V_{1}$ should be the position of unstable equilibrium. The point $V=V_{2}$ can be the point of either stable, or unstable equilibrium. We met with similar dichotomy while studying travelling waves with finite support in Josephson transmission line kogan4 ). The stable equilibrium waves we called shocks, and the unstable equilibrium - kinks. There are several important distinctions between these two types of waves.
Unstable equilibrium at both at the point $V=V_{1}$ and $V=V_{2}$ means that differentiating the r.h.s. of (45) we recover equation (12). The kink, similar to what was in the lossless case, is supersonic both with respect to the region before the kink and the region behind the kink. Since the particle stops at the unstable equilibrium point, for the kink to exist, the fine tuning is necessary - the parameters $V_{1}$, $V_{2}$ and $R$ should satisfy definite relation.
Stable equilibrium at the point $V=V_{2}$ means
$U^{2}L\left(\frac{dQ}{dV}\right)_{V=V_{2}}-1<0.$ (50)
Inequalities (50) and (12) (the latter only for $V=V_{1}$) reflect the well-known in the nonlinear waves theory fact: the shock speed is larger than the sound speed in the region before the shock but smaller than the sound speed in the region behind the shock whitham .
To clarify the distinction between the shocks and the kinks let us bring together the formulas for the travelling wave speed and the sound speed for the cubic case
$\frac{1}{LC_{0}U^{2}}=\alpha+\beta\left(V_{1}+V_{2}\right)+\gamma\left(V_{1}^{2}+V_{1}V_{2}+V_{2}^{2}\right),$ (51a)
$\frac{1}{LC_{0}u^{2}(V_{1,2})}=\alpha+2\beta V_{1,2}+3\gamma V_{1,2}^{2}.$ (51b)
We realize that for $\beta>0$ and $\gamma>0$, for $V_{1}$ and $V_{2}$ having the same sign the travelling wave is always the shock. For $V_{1}$ and $V_{2}$ having opposite signs, we can have both the kink and the shock, depending upon the parameters.
To study quantitatively dissipative kinks and shocks we will use the approach of the previous Section. After repeating exactly what was done there we obtain modified Eq. (10)
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+\frac{UL}{R}\frac{dE}{dV}=U^{2}L\frac{dQ(V)}{dV}-1.$ (52)
The boundary conditions (11) lead to the boundary conditions (15) for $E(V)$, like it was in the lossless case. Equation (52) in the vicinity of $V_{1,2}$ has the form
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+\frac{UL}{R}\frac{dE}{dV}=\text{const},$ (53)
which contains additional term in the l.h.s. with respect to (16), but we still obtain the condition (17). We can also repeat verbatim the part of Section II.2 leading to Eqs. (21) and (22).
In the cubic case Eq. (52) becomes
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+\frac{UL}{R}\frac{dE}{dV}=U^{2}L(\alpha+2\beta V+3\gamma V^{2})-1.$ (54)
We regain Eq. (23), and there exists kogan4 a simple polynomial solution of (54) in the form (24), which we repeat here
$E(V)=\psi(V-V_{1})(V-V_{2}),$ (55)
where $\psi$ is a constant (positive for positive $V_{1}$). Substituting (55) into (54) and equating the coefficients before the same powers of $V$ we obtain after simple algebra the generalization of (II.2)
$\psi^{2}$ $=6U^{2}LC_{0}\gamma,$ (56a)
$V_{1}+V_{2}$ $=-\frac{2\beta}{3\gamma}+\frac{4UL}{\psi R}.$ (56b)
Note that Eqs. (25a), (35a) and (56a) are all identical, hence all the waves we consider in this paper can exist only if $\gamma>0$. Note also that by integrating (54) we can obtain Abel equation of the second kind, and (56b) is the condition of the solvability of the equation polyanin .
Equation (55) shows that in the lossy line the kinks are of the same shape as in the lossless line (given by (27)). And the the shocks are of the same shape also. However, the value of $V_{2}$ (for a given $V_{1}$) depends upon $R$ (Eq. (56b) instead of (25b)) and so does the kink (the shock) speed, in spite of the validity in all the cases of Eq. (23).
The boundary in the phase space between the kinks and the shocks can be easily obtained from (III) in the case when $\beta=0$. Equating the speeds we obtain that this boundary corresponds to
$V_{1}+2V_{2}=0$ (57)
(very much similar to what we had for the Josephson transmission line kogan4 ).
We need to emphasise the different meaning of the condition (56b) for the kinks and the shocks. For the kinks, (56b) can be understood as the constraint on $V_{2}$ (for a given value of $V_{1}$ and ohmic resistance $R$). And it happens that the kink satisfying such constraint is described by the elementary function. For the shock, (56b) is just the condition for it being described by an elementary function. But, of course, the shock with any other value of $V_{2}$ (satisfying (48) also exists.
Let us assume that $V_{1}$ and $V_{2}$ are so close to each other, that we can keep in the r.h.s. of (18) only the first two terms, thus simplifying (54) to
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+\frac{UL}{R}\frac{dE}{dV}=U^{2}LC_{0}(\alpha+2\beta V)-1.$ (58)
The solution of (58) satisfying the boundary conditions (15) and (29) is
$E=\chi\left(V-V_{2}\right)\left[\left(V-V_{2}\right)^{1/2}-\left(\Delta V\right)^{1/2}\right],$ (59)
where $\Delta V\equiv V_{1}-V_{2}$ (we assumed for the sake of definiteness that $V_{1},V_{2}>0$). Substituting (59) into (58) we obtain
$\chi^{2}$ $=8U^{2}LC_{0}\beta$ (60a)
$\frac{UL}{R}$ $=\frac{5\chi}{24}\left(\Delta V\right)^{1/2}.$ (60b)
The function $E(V)$ being found, we can find the function $V(Z)$ by solving equation
$\frac{dV}{dX}=\chi\left(V-V_{2}\right)\left[\left(V-V_{2}\right)^{1/2}-\left(\Delta V\right)^{1/2}\right].$ (61)
The solution is
$V(Z)=V_{2}+\frac{\Delta V}{\left[\exp\left(\chi Z/2\right)+1\right]^{2}}.$ (62)
According to what was said after Eq. (III), Eq. (62) describes the shock. And we have to repeat here what was said after Eq. (III). Equation (60b) is just the condition for the shock to be expressed by an elementary function. The shock exists for any other value of $V_{2}$.
IV Conclusions
We considered finite support travelling waves of the voltage and the current in the discrete lossless transmission line with nonlinear capacitors. We have shown that these waves are of two kinds: the kinks, when the voltage has different asymptotic values at both sides of the transmission line, and the solitons (solitary waves), when the voltage has the same asymptotic value at both sides of the transmission line. We also studied lossy transmission lines. The travelling waves with finite support in such systems are the dissipative kinks and the shocks. We calculated the speeds and the profiles both of the dissipative kinks and of the shocks.
Acknowledgements.
I am grateful to B. Malomed for the discussion. The paper was finalized during the Author’ participation in Workshop on Twistronics and Moire Materials: Bridging Theory and Experiments, ICTP, Trieste. The Author is grateful to the Center for the hospitality.