### v5

# On the kinks, the solitons and the shocks in discrete nonlinear transmission line

Abstract
We studied lossless discrete transmission lines constructed from ideal linear inductors and nonlinear capacitors. The travelling waves with finite support in such systems are the kinks and the solitons, which speeds and profiles were calculated. We also studied lossy transmission lines. The travelling waves with finite support in such systems are the dissipative kinks and the shocks, which speeds and profiles were also calculated.
I Introduction
The nonlinear electrical transmission lines are of much interest both due to their applications, and as the laboratories to study nonlinear waves malomed2 . Among the latter – the travelling waves, especially the kinks and the solitons, attract a lot of attention solitons . The nonlinear waves become especially interesting (and challenging) when one takes into account the discrete nature of the transmission line. The kinks and the solitons in discrete transmission lines were studied in quite a few publications malomed4 ; cameroon ; bogdan .
The field of nonlinear waves in discrete transmission lines is closely connected with Fermi-Pasta-Ulam-Tsingou (FPUT) problem fermi . In particular, dynamics of lattice kinks and solitons were studied in a large number of publications kevrekidis ; vainchtein ; vainchtein2 ; kamchatnov
We studied previously the kinks and the solitons in discrete Josephson transmission line (JTL) JTL kogan2 ; kogan3 . There we formulated the quasi-static approximation, which allows to reduce the discrete equations to the continuous ones. In the present short note we would like to apply this approximation to different kind of discrete transmission lines – with nonlinear capacitors. Such transmission lines in the continuum approximation were studied in the classical papers by R. Landauer landauer1 ; landauer2 ; landauer .
II The kinks and the solitons
II.1 The travelling waves
The discrete transmission line is constructed from the identical inductors and nonlinear capacitors is shown on Fig. 1.
Figure 1: Discrete transmission line.
We take the voltages $V_{n}$ and the currents which pass though the inductors $I_{n}$ as the dynamical variables. The circuit equations (Kirchhoff laws) are
$\frac{dQ_{n}}{dt}$ $=I^{c}_{n},$ (1a)
$L\frac{dI_{n,n+1}}{dt}$ $=V_{n}-V_{n+1},$ (1b)
$I^{c}_{n}$ $=I_{n-1,n}-I_{n,n+1}.$ (1c)
where $Q_{n}=Q(V_{n})$ and the nonlinear nature of the transmission line is determined by the (monotonically increasing) function $Q=Q(V)$. We can exclude $I_{n,n+1}$ from (1b) and (1c) and present these two equations a single one
$L\frac{dI^{c}_{n}}{dt}=V_{n+1}-2V_{n}+V_{n-1}.$ (2)
From (1a) and (2) we can obtain closed wave equation for the voltage $V$
$L\frac{d^{2}Q_{n}}{dt^{2}}=V_{n+1}-2V_{n}+V_{n-1}.$ (3)
Let us write down Eq. (3) in the quasi-continuum approximation, which was proposed in our previous publications kogan2 ; kogan3 . The approximation consists in treating $n$ as the continuous variable $Z$, expanding the r.h.s. of the equation (3) in Taylor series with respect to $Z$ and keeping only the first two terms of the expansion, thus obtaining the equation in the form
$L\frac{\partial^{2}Q(V)}{\partial t^{2}}=\frac{\partial^{2}V}{\partial Z^{2}}+\frac{1}{12}\frac{\partial^{4}V}{\partial Z^{4}}.$ (4)
In this paper we are interested in the travelling wave solutions, for which the dependence of all the variables upon $Z$ and $t$ is of the form
$X(Z,t)=X(Z-Ut),$ (5)
where $U$ is the speed of the travelling wave (the distance is measured in the units of the transmission line spacial period $\Lambda$). This from (4) we obtain an ordinary differential equation
$U^{2}L\frac{d^{2}Q(V)}{dZ^{2}}=\frac{d^{2}V}{dZ^{2}}+\frac{1}{12}\frac{d^{4}V}{dZ^{4}}.$ (6)
Performing an integration we get
$U^{2}L\frac{dQ}{dZ}=\frac{dV}{dZ}+\frac{1}{12}\frac{d^{3}V}{dZ^{3}}.$ (7)
One way to solve Eq. (7) is to perform the second integration to obtain
$\frac{1}{12}\frac{d^{2}V}{dZ^{2}}=U^{2}LQ(V)-V+F,$ (8)
where $F$ is the constant of integration.
However, in this Section we’ll use another method of solution of Eq. (7). After presenting $dQ/dZ$ as $(dQ/dV)(dV/dZ)$, we are motivated to introduce the new dependent variable
$E=\frac{dV}{dZ}.$ (9)
Considering $V$ as the independent variable, taking into account that
$\frac{d^{2}V}{dZ^{2}}$ $=\frac{1}{2}\frac{dE^{2}}{dV},$ (10a)
$\frac{d^{3}V}{dZ^{3}}$ $=\frac{1}{2}E\frac{d^{2}E^{2}}{dV^{2}}$ (10b)
and dividing both parts by the common multiplier $E$ in both parts, we may reduce the order of (7) and write it down as
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}=U^{2}L\frac{dQ(V)}{dV}-1.$ (11)
Consider the ”sound” waves that is small amplitude, smooth waves propagating on the homogeneous backgrounds $V$. These waves are described by Eq. (11), with the l.h.s. being put equal to zero. The speed of the sound is given by the equation
$\frac{1}{u^{2}(V)}=L\frac{dQ}{dV}.$ (12)
Thus Eq. (11) can be presented as
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}=\frac{U^{2}}{u^{2}(V)}-1.$ (13)
Note that in the framework of the quasi-continuum approximation $E$ is just the voltage on the inductor. Thus the physical meaning of Eq. (13) is simple. Its solution gives the voltage on the inductor as the function of the voltage on the capacitor.
A simple mechanical interpretation of Eq. (13) may be helpful. If we call $V$ the ”time”, the equation can be considered as describing motion of the fictitious Newtonian particle, $E^{2}$ being its coordinate and $1/24$ it’s mass, under the influence of the time-dependent force, given by the r.h.s. of the equation.
We’ll consider travelling waves with finite support, that is impose upon $V(Z)$ the boundary conditions
$\lim_{Z\to-\infty}V(Z)=V_{1},\hskip 28.45274pt\lim_{Z\to+\infty}V(Z)=V_{2}.$ (14)
The value of $V_{2}$ may be different from $V_{1}$, in which case we’ll talk about the kink, or equal to $V_{1}$, in which case we’ll talk about the soliton. The quantity $E^{2}(V)$ obviously has minimum (equal to zero) at $V=V_{1,2}$. Hence from (13) follows
$U^{2}>u^{2}\left(V_{1,2}\right),$ (15)
that is the travelling waves are supersonic kogan2 ; kogan3 .
The boundary conditions (14) obviously give
$E\left(V_{1,2}\right)=0.$ (16)
Also, from (10a) and (14) follows that in the vicinity of $V_{1,2}$
$E^{2}\sim\left(V-V_{1,2}\right)^{2}.$ (17)
The conditions (16) and (17) mean that the particle is at rest at the origin at ”time” equal to $V_{1}$ and then again at ”time” equal to $V_{2}$.
II.2 The kinks
With four boundary conditions specifying both the position and the velocity of the particle at $V=V_{1}$ and $V=V_{2}$, of the three parameters $V_{1},V_{2},U^{2}$ only one (say $V_{1}$) can be considered as a free parameter, the other two are connected with it. One such constraint we could obtain at much earlier stage. By integrating (7) with respect to $V$ from $V_{1}$ to $V_{2}$ and taking into account the boundary conditions we obtain the speed of the kink
$\frac{1}{U^{2}}=L\frac{Q_{2}-Q_{1}}{V_{2}-V_{1}}.$ (18)
where $Q_{i}=Q(V_{i})$.
If $Q(V)$ is a polynomial function of $V$, from (17) follows that the kink can be presented as
$E^{2}(V)=P_{1}(V)(V-V_{1})^{2}(V-V_{2})^{2},$ (19)
where $P_{1}(V)$ is some polynomial function.
If we approximate $Q(V)$ between $V_{1}$ and $V_{2}$ by the third degree polynomial
$Q(V)=\text{const}+C_{0}\left(\alpha V+\beta V^{2}+\gamma V^{3}\right)$ (20)
(for brevity we’ll call Eq. (20) the cubic case), then
$\frac{1}{u^{2}(V)}$ $=LC_{0}\left(\alpha+2\beta V+3\gamma V^{2}\right),$ (21a)
$\frac{1}{U^{2}}$ $=LC_{0}\left[\alpha+\beta\left(V_{1}+V_{2}\right)+\gamma\left(V_{1}^{2}+V_{1}V_{2}+V_{2}^{2}\right)\right].$ (21b)
We also can specify (19) to
$E(V)=\psi(V-V_{1})(V-V_{2}),$ (22)
where $\psi$ is a constant (positive, if $V_{1}>V_{2}$, and negative in the opposite case). Substituting (22) into (13) and using Vieta’s formula we obtain after simple algebra
$\psi^{2}$ $=6U^{2}LC_{0}\gamma,$ (23a)
$V_{1}+V_{2}$ $=-\frac{2\beta}{3\gamma}.$ (23b)
Thus the necessary condition for the existence of the kink is the inequality $\gamma>0$, that is the speed of sound $u^{2}(V)$ should be a convex function of $V$.
To find $V(Z)$ we have to solve Eq. (9). The solution is
$V(Z)=\frac{V_{1}+V_{2}}{2}-\frac{V_{1}-V_{2}}{2}\tanh\left[\frac{V_{1}-V_{2}}{2}\psi Z\right].$ (24)
After finding $V(Z)$ we can from (1b) obtain the current
$I(Z)=-\frac{\left(V_{1}-V_{2}\right)}{2UL}\tanh\left[\frac{V_{1}-V_{2}}{2}\psi Z\right].$ (25)
II.3 The solitons
To specify the soliton we need to introduce an additional parameter, and as such we’ll take the extremal value of the voltage $V_{3}$. That is $V$ changes between $V_{1}$ and $V_{3}$, and $|V_{3}-V_{1}|$ we’ll call the amplitude of the soliton. Because $V_{3}$ is the extremal value of $V$ we have
$\left(\frac{dV}{dZ}\right)_{V=V_{3}}=0,\hskip 28.45274pt\left(\frac{d^{2}V}{dZ^{2}}\right)_{V=V_{3}}\neq 0.$ (26)
Hence from (9) and (10a) follows that in the vicinity of $V_{3}$
$E^{2}\sim\left(V-V_{3}\right).$ (27)
Recalling our mechanical analogy we see that the soliton corresponds to the motion of the fictitious particle with zero initial coordinate and velocity at $V=V_{1}$ and zero coordinate (but not velocity) at $V=V_{3}$. Mechanical interpretation of (11) in the case of the soliton leads to the situation which doesn’t normally occur in mechanics. The boundary condition $V_{2}=V_{1}$ means that the ”time” $V$ flows along the contour from $V_{1}$ to $V_{3}$ and then back from $V_{3}$ to $V_{1}$. This is why $E(V)$ for the soliton will later turn out to be a double-valued function.
With the position and the velocity of the particle at $V=V_{1}$ and its position at $V=V_{3}$ being specified, of the three parameters $V_{1},V_{3},U^{2}$ only two (say $V_{1}$ and $V_{3}$) can be considered as free parameters. Multiplying both sides of (7) by $V_{3}-V$, integrating the equation with respect to $V$ from $V_{1}$ to $V_{3}$ and taking into account the boundary conditions we obtain the speed of the soliton
[…]
If $Q(V)$ is a polynomial function of $V$, from (17) and (27) follows that the soliton can be presented as
$E^{2}(V)=P_{2}(V)(V-V_{1})^{2}(V-V_{3}),$ (29)
where $P_{2}(V)$ is some polynomial function.
In the cubic case, calculating the integral in (28) we obtain
$\frac{1}{U^{2}}=LC_{0}\left[\alpha+\frac{2\beta}{3}\left(2V_{1}+V_{3}\right)+\frac{\gamma}{2}\left(3V_{1}^{2}+2V_{1}V_{3}+V_{3}^{2}\right)\right].$
We also can specify (29) to
$E^{2}(V)=\psi(V-V_{1})^{2}(V-V_{3})(V-V_{4}),$ (31)
where $\psi$ is a constant. Substituting (31) into (13) and using Vieta’s formula we obtain after simple algebra
$\psi=6U^{2}LC_{0}\gamma,$ (32a)
$2V_{1}+V_{3}+V_{4}=-\frac{4\beta}{3\gamma}.$ (32b)
To make life easier to ourselves, let us restrict the consideration by the case $|\beta/\gamma|\gg|V_{1}|,|V_{3}|$. Physically the approximation means that $V_{1}$ and $V_{3}$ are close enough to each other, and we can ignore the cubic term in (20). In this case, excluding $V_{4}$ and introducing the amplitude $\phi$ instead of $\psi$ we obtain that (31) and (II.3) are modified to
$E^{2}(V)$ $=\phi(V-V_{1})^{2}(V-V_{3}),$ (33a)
$\frac{\phi}{4}$ $=2U^{2}LC_{0}\beta.$ (33b)
From (II.3) we see that the soliton exists both for positive and negative $\beta$. For positive $\beta$ we have dark soliton ($V_{3}<V_{1}$), and for negative $\beta$ – bright soliton ($V_{3}>V_{1}$).
To find $V(Z)$ we have to solve Eq. (9). The solution is
$V(Z)=V_{1}+\left(V_{3}-V_{1}\right)\operatorname{sech}^{2}\left(\frac{V_{3}-V_{1}}{2}\phi Z\right).$ (34)
After finding $V(Z)$ we can from (1b) obtain the current
$I(Z)=\frac{V_{3}-V_{1}}{2UL}\operatorname{sech}^{2}\left(\frac{V_{3}-V_{1}}{2}\phi Z\right).$ (35)
II.4 The quasi-continuum approximation
We would like to discuss an opportunity (and motivation) to go beyond the limits of the quasi-continuum approximation used in the paper. To do it, one may keep the third term in the Taylor expansion of the discrete second order derivative in the r.h.s. of (2) kogan2 ; kogan3
$V_{n+1}-2V_{n}+V_{n-1}=\frac{\partial^{2}V}{\partial Z^{2}}+\frac{1}{12}\frac{\partial^{4}V}{\partial Z^{4}}+\frac{1}{360}\frac{d^{6}V}{dZ^{6}}.$ (36)
Thus we obtain instead of (7) the equation
$U^{2}L\frac{dQ}{dZ}=\frac{dV}{dZ}+\frac{1}{12}\frac{d^{3}V}{dZ^{3}}+\frac{1}{360}\frac{d^{5}V}{dZ^{5}}.$ (37)
Looking at the results of the previous Sections we understand that the quasi-continuum approximation can be justified if $\gamma(V_{1}-V_{2})^{2}/\alpha\ll 1$ and $\beta\left(V_{1}-V_{3}\right)/\alpha\ll 1$.
The small numerical coefficient before the sixth order derivative in the r.h.s. of (37) means that the small parameters shouldn’t be actually very small. With all that, one can expect that Eq. (37) will allow to study the cases of stronger nonlinearity than Eq. (7).
III The dissipative kinks and shocks
Let us ask ourselves what happens with the kinks when the transmission line becomes lossy? To answer this question let us include in the system the resistors $R$ shunting the inductors (see Fig. 2).
Figure 2: Discrete transmission line with shunted inductors.
In this case the circuit equations are modified to
$L\frac{dI_{n,n+1}}{dt}$ $=\left(1+R_{C}C\frac{d}{dt}\right)\left(V_{n}-V_{n+1}\right),$ (38a)
$I^{c}_{n}$ $=\left(1+\frac{L}{R_{L}}\frac{d}{dt}\right)\left(I_{n-1,n}-I_{n,n+1}\right)$ (38b)
(Eq. (1a) remains as it was). Hence (3) is modified to
$L\frac{d^{2}Q_{n}}{dt^{2}}=\left(1+\frac{L}{R_{L}}\frac{d}{dt}\right)\left(1+R_{C}C\frac{d}{dt}\right)$
$\left(V_{n+1}-2V_{n}+V_{n-1}\right),$ (39)
and (4) is modified due to the presence of ohmic resistors to
$L\frac{\partial^{2}Q(V)}{\partial t^{2}}=\frac{\partial^{2}V}{\partial Z^{2}}+\tau_{R}\frac{\partial^{3}V}{\partial t\partial^{2}Z}+\frac{1}{12}\frac{\partial^{4}V}{\partial Z^{4}},$ (40)
where
$\tau_{R}=\frac{L}{R_{L}}+R_{C}C.$ (41)
In deriving (40) we assumed that both the influence of the higher order derivatives with respect to $Z$ and the dissipation are small, so we ignored the terms which which correspond to the products of these effects.
For the travelling wave we obtain from (40) after double integration the generalization of (8)
$\frac{1}{12}\frac{d^{2}V}{dZ^{2}}+U\tau_{R}\frac{dV}{dZ}=U^{2}LQ(V)-V+F.$ (42)
Equation (42) is Newton equation, describing motion with friction of the fictitious particle ($V$ being the coordinate of the particle and $Z$ the ”time”) in the force field given by the r.h.s. of the equation. We again impose on $V(Z)$ the boundary conditions (14). Hence the motion starts at $Z=-\infty$ at $V=V_{1}$ and ends at $Z=+\infty$ at $V=V_{2}$ (obviously that both $V_{1}$ and $V_{2}$ are the equilibrium positions). Equating the r.h.s. of (42) to zero at $V=V_{1}$ and $V=V_{2}$ we recover Eq. (18) for the speed of the wave.
It will be convenient to present Eq. (42) as
$\frac{1}{12}\frac{d^{2}V}{dZ^{2}}+U\tau_{R}\frac{dV}{dZ}=-\frac{d\Pi(V)}{dV},$ (43)
where the ”potential energy” $\Pi(V)$ is defined by the equation (which took into account the boundary conditions)
$\frac{d\Pi(V)}{dV}=-\frac{V_{2}-V_{1}}{Q_{2}-Q_{1}}Q(V)+V+\frac{V_{2}Q_{1}-V_{1}Q_{2}}{Q_{2}-Q_{1}}.$ (44)
The constraint for the travelling wave asymptotic boundary conditions is the inequality
$\Pi(V_{1})>\Pi(V_{2}).$ (45)
Inequality (45) can be explicitly presented as
[…]
For $\gamma>0,\beta\geq 0$, the condition (46) means just $|V_{1}|>|V_{2}|$.
The point $V=V_{1}$ should be the position of unstable equilibrium. The point $V=V_{2}$ can be the point of either stable, or unstable equilibrium. (We met with similar dichotomy while studying travelling waves with finite support in Josephson transmission line kogan4 ). The waves corresponding to the stable equilibrium at the point $V=V_{2}$ we called the shocks, to the unstable equilibrium – the kinks.
Differentiating (44) we obtain the r.h.s. of (49) (see below)
$-\frac{d^{2}\Pi(V)}{dV^{2}}=\frac{U^{2}}{u^{2}(V)}-1;$ (47)
the point $V=V_{i}$ being the point of the stable/unstable equilibrium corresponds to the travelling wave being subsonic/supersonic with respect to the background. Stable equilibrium at the point $V=V_{2}$ (together with unstable equilibrium at the point $V=V_{2}$) reflect the well-known in the nonlinear waves theory fact: the shock speed is larger than the sound speed in the region before the shock but smaller than the sound speed in the region behind the shock whitham .
Let us return to (II.2) giving the sound and the kink/shock speeds for the cubic case. The kinks/shocks phase diagram in the space $V_{1},V_{2}$ can be obtained by comparing $U^{2}$ and $u^{2}(V_{2})$. Consider the case $\gamma>0,\beta\geq 0$. If $V_{1}$ and $V_{2}$ have the same sign, the travelling wave is always the shock. In the opposite case, the wave can be either the kink or the shock, depending upon $V_{1}$ and $V_{2}$. The boundary in the phase space between the kinks and the shocks is especially simple for $\beta=0$
$V_{1}+2V_{2}=0$ (48)
(very much similar to what we had for the Josephson transmission line kogan4 ).
The most important distinction of the dissipative kinks from the shocks: since the particle stops at the unstable equilibrium point, for the kink to exist, the fine tuning is necessary - the parameters $V_{1}$, $V_{2}$ and $R$ should satisfy definite relation (see below). The shocks don’t need such find tuning – they are generic.
To study quantitatively dissipative kinks and shocks we will use the approach of the previous Section. After repeating exactly what was done there we obtain modified Eq. (11)
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+U\tau_{R}\frac{dE}{dV}=\frac{U^{2}}{u^{2}(V)}-1.$ (49)
At this stage we can repeat verbatim the part of Section II.2 leading to (22), hence the dissipative kinks and the shocks are of the same shape (given by (24)) as the kinks in the lossless line.
However, Eq. (II.2) is generalized to
$\psi^{2}$ $=6U^{2}LC_{0}\gamma,$ (50a)
$V_{1}+V_{2}$ $=-\frac{2\beta}{3\gamma}+\frac{4U\tau_{R}}{\psi}.$ (50b)
hence the value of $V_{2}$ (for a given $V_{1}$) depends upon $R$ (Eq. (50b) instead of (23b)) and so does the kink (the shock) speed and amplitude, in spite of the validity in all the cases of Eq. (21b).
We need to emphasise the different meaning of the condition (50b) for the kinks and the shocks. For the kinks, (50b) is the constraint on $V_{1},V_{2},R$. And it happens that the kink satisfying such constraint is described by the elementary function. For the shock, (50b) is just the condition for it being described by an elementary function. But, of course, the shock with any other values of $V_{1},V_{2},R$ also exists (provided that $V_{1},V_{2}$ satisfy the constraint (45)). Note also that by integrating (49) we can obtain Abel equation of the second kind, and (50b) is the condition of the solvability of the equation polyanin .
Let us assume that $V_{1}$ and $V_{2}$ are so close to each other, that we can keep in the r.h.s. of (20) only the first two terms, thus simplifying (49) to
$\frac{1}{24}\frac{d^{2}E^{2}}{dV^{2}}+U\tau_{R}\frac{dE}{dV}=U^{2}LC_{0}(\alpha+2\beta V)-1$ (51)
(we assume for the sake of definiteness that $\beta>0$ and $V_{1},V_{2}>0$). The elementary solution of (51) satisfying the boundary conditions (16) is of the form kogan4
$E=\chi\left(V-V_{2}\right)\left[\left(V-V_{2}\right)^{1/2}-\left(V_{1}-V_{2}\right)^{1/2}\right].$ (52)
Substituting (52) into (51) we obtain
$\chi^{2}$ $=8U^{2}LC_{0}\beta$ (53a)
$U\tau_{R}$ $=\frac{5\chi}{24}\left(V_{1}-V_{2}\right)^{1/2}.$ (53b)
Similar to Eq. (50b) in the case of shocks, Eq. (53b) is just the condition for the shock to be expressed by an elementary function. The shock exists for any other value of $\Delta V$.
The function $E(V)$ being found, we can find the function $V(Z)$ by solving Eq. (9). The solution is
$V_{2}(Z)=V_{2}+\frac{V_{1}-V_{2}}{\left[\exp\left(\chi Z/2\right)+1\right]^{2}}.$ (54)
IV Conclusions
We considered finite support travelling waves of the voltage and the current in the discrete lossless transmission line with nonlinear capacitors. We have shown that these waves are of two kinds: the kinks, when the voltage has different asymptotic values at both sides of the transmission line, and the solitons (solitary waves), when the voltage has the same asymptotic value at both sides of the transmission line. We also studied lossy transmission lines. The travelling waves with finite support in such systems are the dissipative kinks and the shocks. We calculated the speeds and the profiles both of the dissipative kinks and of the shocks.
Acknowledgements.
I am grateful to A. M. Kamchatnov, B. Malomed and A. Vainchtein for the discussion. The paper was finalized during the Author’ participation in Workshop on Twistronics and Moire Materials: Bridging Theory and Experiments, ICTP, Trieste. The Author is grateful to the Center for the hospitality.